OpenStudy (anonymous):

Find the value of the discriminant describe the number and type of roots for the equation. Then find the exact solution of the following quadratic equation by using the Quadratic Formula. 4x^2 = x + 7

hero (hero):

Re-write the equation in standard form: \[4x^2 - x - 7 = 0\] Now use the quadratic formula: \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] \[x = \frac{-(-1) \pm \sqrt{(-1)^2 - 4(4)(-7)}}{2(4)}\]